OpenStudy (anonymous):

Graphing Linear Equation Help (Medal) I have graphed a linear equation where it is obvious the lines intersect, but not on an exact point. How do I find the solution?

OpenStudy (anonymous):

example

jimthompson5910 (jim_thompson5910):

we have these two equations y = 3x x+2y = 4

jimthompson5910 (jim_thompson5910):

Start with the equation x + 2y = 4 and replace y with 3x (since y = 3x) to get... x+2y = 4 x + 2(3x) = 4 ... replace y with 3x x + 6x = 4 do you know how to solve from here?